OpenStudy (anonymous):

A semicircular window of radius 12 inches is to be laminated with a sunblock coating that costs $0.80 per square inch to apply. What is the total cost of coating the window, to the nearest cent? (Use p = 3.14. Hint: Area of Circle: A = p r²

OpenStudy (anonymous):

Okay, we have the following as given: We have a semicircle with R = 12 in so: \[A = 1/2 \pi r^2 = 72\pi\] Now the other half of the circle also has the same Radius = 12 in so it'll have the same area, \[A = 1/2 \pi r^2 = 72\pi\] and so If 72 pi of the circle ---> costs 0.80 then 72 pi + 72 pi (full circle) ----> x cross multiply and you'll get: x = 1.6 $ It'll cost 1.6 $ for total coating ^_^ correct me if I'm wrong

OpenStudy (anonymous):

A.) $180.86 B.) $361.73 C.) $60.29 D.) $1,808.60

OpenStudy (anonymous):

lol, I got none, so I guess I'm wrong ^^"

OpenStudy (anonymous):

That's okay.. I have another for ya if you wanna try

OpenStudy (anonymous):

sure ^_^